MATH 367, fall 98 QUIZ 1, Sept 16

NAME:

1. Consider the words of 6 characters that can be formed with the characters a, b, c, d, e, f, 1, 2, 3, 4.

(a) How many words can be formed if the second and last character must be a number, and repetitions are not allowed.

(b) Same as in 1a but repetitions are allowed.

(c) How many words can be formed if no repetitions are allowed and there must be 4 letters and 2 numbers.

(d) How many words can be formed if no repetitions are allowed and there must be at least 4 letters.

2. What is the probability that a poker hand (5 cards) contains 2 Aces?

MATH 367, fall 98 QUIZ 1, Sept 16 Take Home portion: Due Sept 24

NAME:

1. A store offers 3 kinds of cans of sodas. 10 thirsty people come to the store to buy one soda each. How many different selections of sodas are possible?

2. You have 10 identical Red balls and 12 identical Yellow balls. In how many ways can you put them in a row so that you have exactly 2 groups of Red balls, and the first and last balls in the row must be Yellow?